A monkey is decending from the branch of a tree with constant acceleration. If the breaking strength is 75% of the weight of the monkey, the minimum acceleration with which monkey can slide down without breaking the branch is
    (a)    g
    (b)    3g/4
    (c)    g/4
    (d)    g/2

Grade:11

4 Answers

Navjyot Kalra
10 years ago

(c)
Let T be the tension in the branch of a tree when monkey is descending with acceleration a. Then mg - T = ma; and T = 75% of weight of monkey
?    ma = mg - (75/100)  mg= (1/4)  mg
or a =g/4.

(c)
Let T be the tension in the branch of a tree when monkey is descending with acceleration a. Then mg - T = ma; and T = 75% of weight of monkey
?    ma = mg - (75/100)  mg= (1/4)  mg
or a =g/4.
DR zia
5 years ago
As ma=mg _T ..so T=75%=2/3...
Now as a=g.
So put ma=mg ...mg=mg_2/3..mg cross on both sides ..so 1_0.75 ..as2/3 sequel to 0.75...so ans =0.25 or g/4
Piyush raj
5 years ago
A body kept on a smooth inclined plane inclination 1 in x will remain stationary relative to the inclined plane if the plane is given a horizontal accleraaccel equal to
Danny
5 years ago
Here mg - T = ma
          mg - (75/100)mg = ma
          a = g/4
..................................................................